OpenStudy (anonymous):

Solve for x: 3 over 4 x + 5 over 8 = 4x Write your answer as a fraction in simplest form.

OpenStudy (fefetasprites):

You need to find the Least Common Denominator, and multiply both sides by it. Do you know how to do that?

OpenStudy (anonymous):

\[\frac{ \[\frac{ 3 }{ 4x+5 }\] }{ 8=4x }\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

 like that